Can someone please help me with me with this problem? How many gallons will fit in the box? 1ft cubed = 7.5 gallons. The three given sides of the box are 5ft, 10ft, and 6ft. Thank you!
Answer by mananth(16946)

1ft cubed = 7.5 gallons.

The three given sides of the box are 5ft, 10ft, and 6ft
Volume = 5*10*6 = 300 cu.ft
1 cu.ft = 7.5 gallons
300 cu.ft = 7.5*300
= 2250 gallons